# Think Then Embed: Generative Context Improves Multimodal Embedding

### v2

@@ 7 Ablations
w/ SFT-ed
Reasoner
Per Meta-Task Score Average Score
@@ 7.1 Understanding the role of TTE Reasoner and ECR
A natural question is whether the ECR text alone is sufficient for retrieval, potentially eliminating the need to train a dedicated encoder. However, our experiments on MMEB show that this is not the case. Specifically, when we use the same ECR and apply a text-to-text (T2T) retrieval baseline—by directly encoding the ECR with a strong off-the-shelf text encoder, Jina-V3 (Sturua et al., 2024) – the performance is significantly worse.
Another concern is that imperfect ECRs might introduce noise that propagates into the Embedder, degrading retrieval performance. Interestingly, we find that our TTE Embedder is robust to such noise: it learns to extract useful information from the ECR without blindly relying on it. In fact, although ECRs alone do not yield high retrieval precision, they provide valuable signals that our Embedder can leverage. We observe similar trends across all task types. These findings indicate that our Embedder is robust to noisy ECRs, and retrieval performance can benefit even from imperfect reasoning traces.
@@ 7.3 Embedding Head Design
To better exploit representations from a frozen MLLM, we shift focus to earlier layers (e.g., the 8th-to-last) and equip them with more expressive multi-layer embedding heads. Specifically, we compare a QFormer-style head with a self-initialized multi-head self-attention (MHSA) module. Empirically, the latter achieves superior performance compared to the QFormer-style head, which requires training from scratch. When self-initialized with the last $\text{n\_layers}=8$, the MHSA head reaches performance on par with a separately fine-tuned embedder (TTEs).

### v4

@@ 7 Ablations
w/ SFT-ed Reasoner Per Meta-Task Score Average Score
@@ 7.1 Understanding the role of TTE Reasoner and ECR
A natural question is whether the ECR text alone is sufficient for retrieval, potentially eliminating the need to train a dedicated encoder. However, our experiments on MMEB show that this is not the case. Specifically, when we use the same ECR and apply a text-to-text (T2T) retrieval baseline—by directly encoding the ECR with a strong off-the-shelf text encoder, Jina-V3 (jinav3) – the performance is significantly worse.
Another concern is that imperfect ECRs might introduce noise that propagates into the Embedder, degrading retrieval performance. Interestingly, we find that our TTE Embedder is robust to such noise: it learns to extract useful information from the ECR without blindly relying on it. In fact, although ECRs alone do not yield high retrieval precision, they provide valuable signals that our Embedder can leverage. We observe similar trends across all task types. These findings indicate that our Embedder is robust to noisy ECRs, and retrieval performance can benefit even from imperfect reasoning traces. Meanwhile, we also note that the observed robustness does not imply that the embedder is invariant to the quality of ECRs. In fact, the robustness range we measure is within the same reasoner: the embedder outperforms directly using the ECRs themselves (e.g., for VQA-style tasks where retrieval does not rely on dense visual embeddings). However, using a different reasoner can still impact the embedder’s performance, as shown in Table 3 and between TTEs and TTEt.
@@ 7.3 Embedding Head Design
To better exploit representations from a frozen MLLM, we shift focus to earlier layers (e.g., the 8th-to-last) and equip them with more expressive multi-layer embedding heads. Specifically, we compare a QFormer-style head with a self-initialized multi-head self-attention (MHSA) module. Empirically, the latter achieves superior performance compared to the QFormer-style head, which requires training from scratch. When self-initialized with the last $\text{n\_layers}=8$, the model reaches performance on par with a separately fine-tuned embedder.